• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 23:59:40 on February19,2002

    Entry number 83217

    keyword=Swing Shift summary

    16:00 Begin shift by changing to 3F kinematics. Previous shift finished 3B kinematics. Hall in controlled access to change calorimeter.

    16:15 L-arm momentum and angle set to 15.47 degrees and 3.362 GeV/c.

    Note: Left Arm Dipole GUI reads M3/M4 Over Temp Fault, Xiaodong and Rusty will look into it.

    16:35 Rusty said that the Fault is a faulty reading. Momentum of the dipole is fine.
    16:44 Hall A is in Beam Permit, and ready for Beam

    16:52 Request Controlled Access to check the pulsar, since we don't have scaler information for the left spectrometer.
    18:25 A fused was blow out for a NIM crate on the Right arm. Pulsar moved to another crate. Note: There are two NIM crates with blown fuses.
    18:38 Called MCC to request Beam Permit. 20 uA unrastered beam requested.

    Run # 2689 is a test run. Run # 2690 is a junk run. Run # 2691 is a cosmic run.

    Begin Kinematics 3F

    18:59-19:14 Run # 2692, Carbon, moff, Veto off, 100 K events
    19:15 Call MCC to mask Target to move to LH2, non. rad.
    19:21 Call MCC to ask for 30 uA of beam with raster on.
    19:25-19:42 Run # 2693, LH2 non rad, moff, Veto off
    19:44 Turn Calorimeter magnet on at 600 A.
    19:45-20:00 Run # 2694, LH2 non rad, mon, Veto off
    20:00 Asked MCC to mask target to move target to LH2 rad and change spectrometer.
    20:07 Left spectrometer moved to 16.34 degrees. Momentum
    setting to 3.118 GeV/c
    20:05 MCC called to report that they will need some time to deliver the beam.
    20:14 Left spectrometer ready for beam.
    21:00 Called MCC to check status of accelerator. See ELOG # 175431. It may be another hour or two before beam.
    22:55 The crew chief from MCC called. Most of the problems have been fixed, except for one RF zone that they will try to bypass. This will cause more beam trips until fixed. They are also trying to clear a 2KW Dump FSD that doesn't want to clear. See ELOG # 175444. They expect to deliver beam in ~ 30 minutes.

    23:59 End shift still waiting for beam.